# Stable Differentiable Modal Synthesis for Learning Nonlinear Dynamics

### v2

@@ 2  NONLINEAR TRANSVERSE STRING VIBRATION
$\mathcal{L}u=\mathcal{F}+\mathcal{F}_{\mathrm{e}}.$
@@ 2.2  Nonlinearity
$\mathcal{V}(\xi)=\Big(\sqrt{1+\xi^{2}}-1\Big)^{2}.$
The potential (2.2) can be approximated by Taylor series as $\frac{1}{4}\xi^{4}$ which corresponds to the model used in our previous work [38]. Further simplification can lead to a Kirchhoff-Carrier model [15, 4], where the nonlinearity is averaged over the length of the string. Compared to the Kirchhoff-Carrier model, which adequately reproduces only the pitch glide effect, the potential (2.2) is capable of capturing other perceptually important effects such as phantom partials.
@@ 2.3  Plucking Excitation
$f_{\mathrm{e}}(t)=\begin{cases}\frac{1}{2}f_{\mathrm{amp}}\big[1-\cos\big(\frac{\pi t}{T_{\mathrm{e}}}\big)\big],\quad&t\in[0,T_{\mathrm{e}}]\\ 0,&\text{otherwise}\end{cases}$
@@ 2.4  Equation Scaling
$\partial_{t}^{2}\hat{u}$ $=\gamma^{2}\partial_{\hat{x}}^{2}\hat{u}-\kappa^{2}\partial_{\hat{x}}^{4}\hat{u}-2\sigma_{0}\partial_{t}\hat{u}+2\hat{\sigma}_{1}\partial_{t}\partial_{\hat{x}}^{2}\hat{u}$+{}
${}+\gamma^{2}\frac{\alpha^{2}-1}{2}\partial_{\hat{x}}\big(\mathcal{V}^{\prime}(\xi)\big)+\delta(\hat{x}-\hat{x}_{\mathrm{e}})\hat{f}_{\mathrm{e}}(t),$ (4)
@@ 3  MODAL DECOMPOSITION
$u(x,t)=\sum\limits_{m=1}^{M}\Phi_{m}(x)q_{m}(t)=\boldsymbol{\Phi}^{\mathrm{T}}(x)\mathbf{q}(t),$
@@
$\ddot{\mathbf{q}}+2\bm{\Sigma}\dot{\mathbf{q}}+\bm{\Omega}^{2}\mathbf{q}=\nu^{2}\mathbf{f}(\mathbf{q})+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t),$
@@
$\begin{dcases}\dot{\mathbf{q}}=\mathbf{p}\\ \dot{\mathbf{p}}=-2\bm{\Sigma}\mathbf{p}-\bm{\Omega}^{2}\mathbf{q}+\nu^{2}\mathbf{f}(\mathbf{q})+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t)\end{dcases}$
@@ 3.1  Spectral Method
To obtain a closed-form expression for the dimensionless nonlinearity $\mathbf{f}(\mathbf{q})\!\!:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$, we employ a spectral method [37, 36] to calculate spatial derivatives $\partial_{x}$ on the grid $x_{l+\nicefrac{{1}}{{2}}}=\frac{1}{M+1}\big(l+\frac{1}{2}\big),\;l=0,\dots,M$. Compared to our previous work [38] that calculated spatial derivatives $\partial_{x}$ over a continuous domain, the spectral method results in a significantly more efficient expression for $\mathbf{f}(\mathbf{q})$ and allows to implement the nonlinear function (2.2) without approximation by Taylor series.
@@
[…]
@@
$\mathbf{f}(\mathbf{q})=-\nabla_{\mathbf{q}}V(\mathbf{q}),\quad V(\mathbf{q})=\frac{1}{M+1}\sum\limits_{l=0}^{M}\mathcal{V}\big(\xi_{l+\nicefrac{{1}}{{2}}}\big).$
@@ 4.1  Quadratisation
where $\epsilon>0$ is a gauge constant. The auxiliary variable $\psi$ is well defined if the potential $V(\mathbf{q})$ (3.1) is non-negative which is the case for the chosen nonlinear function (2.2).
@@
$\begin{dcases}\dot{\mathbf{q}}=\mathbf{p}\\ \dot{\mathbf{p}}=-2\bm{\Sigma}\mathbf{p}-\bm{\Omega}^{2}\mathbf{q}-\nu^{2}\mathbf{g}\psi+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t)\\ \dot{\psi}=\mathbf{g}^{\mathrm{T}}\mathbf{p}\end{dcases}$
@@
Furthermore, we add a control term to the system (4.1) as described by Risse et al. [32] to reduce the drift between numerical values of $\psi$ and $\sqrt{2V(\mathbf{q})+\epsilon}$ in discrete time. This results in modification of the coupling term $\mathbf{g}$:
@@ 4.2  Time Discretisation
$\begin{dcases}\delta_{t+}\mathbf{q}^{n-\nicefrac{{1}}{{2}}}=\mathbf{p}^{n}\\ \begin{aligned} \delta_{t+}\mathbf{p}^{n}=&-2\bm{\Sigma}\mu_{t+}\mathbf{p}^{n}-\bm{\Omega}^{2}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}-{}\\ &{}-\nu^{2}\mathbf{g}^{n}\mu_{t+}\psi^{n}+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}\end{aligned}\\ \delta_{t+}\psi^{n}=(\mathbf{g}^{n})^{\mathrm{T}}\mu_{t+}\mathbf{p}^{n}\end{dcases}$
@@
Numerical energy for the scheme (4.2) is conserved and takes the form:
@@
Defining $\mathbf{q}^{n}\triangleq\mu_{t+}\mathbf{q}^{n-\nicefrac{{1}}{{2}}}$, the scheme (4.2) can be written in an update form $(\mathbf{q}^{n},\mathbf{p}^{n},\psi^{n})\rightarrow(\mathbf{q}^{n+1},\mathbf{p}^{n+1},\psi^{n+1})$:
$\begin{dcases}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}=\mathbf{q}^{n}+\tfrac{k}{2}\mathbf{p}^{n}\\ \begin{aligned} \mathbf{p}^{n+1}&=\Big[\mathbf{I}+k\bm{\Sigma}+\tfrac{k^{2}\nu^{2}}{4}\mathbf{g}^{n}(\mathbf{g}^{n})^{\mathrm{T}}\Big]^{-1}\times{}\\ &{}\times\Big[\Big(\mathbf{I}-k\bm{\Sigma}-\tfrac{k^{2}\nu^{2}}{4}\mathbf{g}^{n}(\mathbf{g}^{n})^{\mathrm{T}}\Big)\mathbf{p}^{n}+{}\\ &{}+k\Big(-\bm{\Omega}^{2}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}-\nu^{2}\mathbf{g}^{n}\psi^{n}+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}\Big)\Big]\end{aligned}\\ \mathbf{q}^{n+1}=\mathbf{q}^{n+\nicefrac{{1}}{{2}}}+\tfrac{k}{2}\mathbf{p}^{n+1}\\ \psi^{n+1}=\psi^{n}+k(\mathbf{g}^{n})^{\mathrm{T}}\tfrac{\mathbf{p}^{n+1}+\mathbf{p}^{n}}{2}\end{dcases}$
The inverse appearing in (4.2) can be easily computed with the Sherman-Morrison formula [35]. Using (3), we obtain an audio output as $w^{n}=\bm{\Phi}^{\mathrm{T}}(x_{\mathrm{o}})\mathbf{q}^{n}$.
@@ 5.1  Neural Ordinary Differential Equations
$\frac{d\mathbf{y}}{dt}=\mathbf{h}_{\theta}(\mathbf{y},t),\quad\mathbf{y}(0)=\mathbf{y}_{0}.$
Here $\mathbf{y}=\mathbf{y}(t)\!\!:\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is an unknown function of time $t$, $\mathbf{y}_{0}\in\mathbb{R}^{K}$ is an initial condition and $\mathbf{h}_{\theta}(\mathbf{y},t)\!\!:\mathbb{R}^{K}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is a neural network where $\theta$ denotes the set of all learnable parameters and $K$ denotes the state dimension. Generally, a simple architecture such as an MLP is chosen for $\mathbf{h}_{\theta}(\mathbf{y},t)$. Chen et al. [6] have showed that the system (5.1) in combination with a numerical solver, labelled as ODENet, can be trained from the observed state data to reproduce the dynamics of a target system for which the theoretical model may be unknown.
Assume a target trajectory $\{\mathbf{y}_{0},\mathbf{y}^{1},\dots,\mathbf{y}^{N-1}\}$ sampled on the time grid $\{t^{n}\}_{n=0}^{N-1}$. Given a predicted trajectory $\{\mathbf{y}_{0},\tilde{\mathbf{y}}^{1},\dots,\tilde{\mathbf{y}}^{N-1}\}$ by a numerical solution to the initial value problem (5.1), i.e., a forward pass of the ODENet, we can construct an objective function $J(\theta)$ such as mean squared error (MSE):
$J(\theta)=\operatorname{MSE}(\tilde{\mathbf{y}}^{n},\mathbf{y}^{n})\triangleq\frac{1}{KN}\sum_{n=0}^{N-1}\left\lVert\tilde{\mathbf{y}}^{n}-\mathbf{y}^{n}\right\rVert_{2}^{2},$
@@ 5.2  Extension for Modal Synthesis
  • set $\bm{\Sigma},\bm{\Omega},\nu,\bm{\Phi}(x_{\mathrm{e}})$ in (5.2) using physical parameters of a target solution;
  • precompute $f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}=f_{\mathrm{e}}\big(t^{n+\nicefrac{{1}}{{2}}}\big),\;n=0,\dots,N-1$ using (2.3) and excitation parameters of a target solution;
  • use the numerical solver (4.2) as in the case of a regular system to produce a predicted trajectory.
The formulation (5.2) has strong implications. The exact expression for linear vibration exploits the periodic, harmonic and lossy nature of a musical system. Thus, we aid the optimisation process by constraining the space of possible solutions and improve interpretability of the model. Furthermore, the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ is memoryless and dimensionless, thus does not depend on physical parameters of a system and external excitation. Theoretically, these parameters can be changed after the training as long as range of displacements $\mathbf{q}$ stays the same as in a training dataset to simulate other configurations of a system. Some limitations exist since the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ implicitly depends on boundary conditions and the number of modes.
@@ 5.3  Gradient Networks
$\mathbf{f}_{\theta}(\mathbf{q})=-\mathbf{W}^{\mathrm{T}}[\bm{\alpha}\odot\sigma(\mathbf{z})],\quad\mathbf{z}=\bm{\beta}\odot\mathbf{W}\mathbf{q}+\mathbf{b},$
@@
We can notice a strong resemblance between the GradNet (5.3) and the spectral method (Section 3.1). The weight matrix $\mathbf{W}$ acts as a transform to a hidden space where we apply the elementwise nonlinear function $\sigma(\mathbf{z})$ before performing an “inverse” transform with the transposed weight matrix $\mathbf{W}^{\mathrm{T}}$. Moreover, the potential (3.1) arising from the spectral method is expressed as a sum of convex ridge functions and the GradNet (5.3) with a monotonically-increasing activation function $\sigma$ and non-negative scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ can universally approximate gradients of such functions [5].
@@ 6  EVALUATION
For the parametrisation of $\mathbf{f}_{\theta}(\mathbf{q})$ we used the GradNet (5.3) with hidden dimension $H=1000$. This number was chosen based on available GPU memory since the validation loss was found to consistently decrease with increasing hidden dimension. However, smaller networks may also be sufficient from the perceptual point of view. A leaky rectified linear unit was used as an activation $\sigma$ which is a monotonically-increasing function with a non-negative antiderivative $\phi$. Kaiming initialisation [12] was used for the initial weights $\mathbf{W}$ and the biases $\mathbf{b}$ were initialised to zero. Logarithms of scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ were used as learnable parameters to enforce non-negativity and were initialised around zero using a normal distribution.
@@
The training loss was the MSE (5.1) taken over the state vector $\mathbf{y}^{\mathrm{T}}=[\mathbf{q}^{\mathrm{T}},\mathbf{p}^{\mathrm{T}}]$, excluding the auxiliary variable $\psi$ arising from the SAV method. This implied that the potential $V_{\theta}(\mathbf{q})$ could be shifted by a constant relative to the target potential (3.1) since dynamics of the system and the training loss would not be affected. Backpropagation was performed using internal operations of the numerical solver (4.2), i.e., the “discretise-then-optimise” method, which is generally a preferred approach due to its gradient accuracy, speed and straightforward implementation [13]. Calculation of the control term $\mathbf{g}_{\mathrm{mod}}$ was excluded from the computational graph. In some cases, it was found to significantly compromise training by steering the potential $V_{\theta}(\mathbf{q})$ to incorrect values of the auxiliary variable $\psi$ and lead to an unexpected increase in the training loss. As only gradient calculation is affected, stability and drift regulation properties of the numerical solver (4.2) still hold during inference. The Adam optimiser [14] was used with default parameters. The training was performed for 2000 epochs using a cluster of four GPUs to parallelise batches of data. The resulting model was chosen based on the lowest validation loss obtained during optimisation.
@@
For training we used a variation of the teacher forcing technique [11] by splitting up a target trajectory into $\mathrm{1}$$\mathrm{ms}$ segments and providing true initial conditions for each segment to the ODENet. Since the numerical solver (4.2) is given as a one step update we had access to both displacement and velocity of the target numerical solution at each time step, and thus initial conditions for each segment. The auxiliary variable was initialised as $\psi_{0}=\sqrt{2V_{\theta}(\mathbf{q}_{0})+\epsilon}$ using initial displacements $\mathbf{q}_{0}$ to produce a consistent initial condition [32]. In addition, the excitation function (2.3) was shifted in time to reflect a new starting point for integration. The main reason for using this technique was to speed up training as the number of integration steps in the numerical solver is significantly reduced. These integration steps cannot be parallelised in time. Moreover, the likelihood of vanishing and exploding gradient problems during optimisation is also reduced by this technique as backpropagation on long time series is avoided [13].

### v3

@@ 2  NONLINEAR TRANSVERSE STRING VIBRATION
$\mathcal{L}u=\mathcal{F}+\mathcal{F}_{\mathrm{e}}.$ (1)
@@ 2.2  Nonlinearity
$\mathcal{V}(\xi)=\Big(\sqrt{1+\xi^{2}}-1\Big)^{2}.$ (2)
The potential (2) can be approximated by Taylor series as $\frac{1}{4}\xi^{4}$ which corresponds to the model used in our previous work [38]. Further simplification can lead to a Kirchhoff-Carrier model [15, 4], where the nonlinearity is averaged over the length of the string. Compared to the Kirchhoff-Carrier model, which adequately reproduces only the pitch glide effect, the potential (2) is capable of capturing other perceptually important effects such as phantom partials.
@@ 2.3  Plucking Excitation
$f_{\mathrm{e}}(t)=\begin{cases}\frac{1}{2}f_{\mathrm{amp}}\big[1-\cos\big(\frac{\pi t}{T_{\mathrm{e}}}\big)\big],\quad&t\in[0,T_{\mathrm{e}}]\\ 0,&\text{otherwise}\end{cases}$ (3)
@@ 2.4  Equation Scaling
$\begin{split}\partial_{t}^{2}\hat{u}&=\gamma^{2}\partial_{\hat{x}}^{2}\hat{u}-\kappa^{2}\partial_{\hat{x}}^{4}\hat{u}-2\sigma_{0}\partial_{t}\hat{u}+2\hat{\sigma}_{1}\partial_{t}\partial_{\hat{x}}^{2}\hat{u}+{}\\ &{}+\gamma^{2}\frac{\alpha^{2}-1}{2}\partial_{\hat{x}}\big(\mathcal{V}^{\prime}(\xi)\big)+\delta(\hat{x}-\hat{x}_{\mathrm{e}})\hat{f}_{\mathrm{e}}(t),\end{split}$ (4)
@@ 3  MODAL DECOMPOSITION
$u(x,t)=\sum\limits_{m=1}^{M}\Phi_{m}(x)q_{m}(t)=\boldsymbol{\Phi}^{\mathrm{T}}(x)\mathbf{q}(t),$ (5)
@@
$\ddot{\mathbf{q}}+2\bm{\Sigma}\dot{\mathbf{q}}+\bm{\Omega}^{2}\mathbf{q}=\nu^{2}\mathbf{f}(\mathbf{q})+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t),$ (6)
@@
$\begin{cases}\dot{\mathbf{q}}=\mathbf{p}\\ \dot{\mathbf{p}}=-2\bm{\Sigma}\mathbf{p}-\bm{\Omega}^{2}\mathbf{q}+\nu^{2}\mathbf{f}(\mathbf{q})+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t)\end{cases}$ (7)
@@ 3.1  Spectral Method
To obtain a closed-form expression for the dimensionless nonlinearity $\mathbf{f}(\mathbf{q})\!\!:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$, we employ a spectral method [37, 36] to calculate spatial derivatives $\partial_{x}$ on the grid $x_{l+\nicefrac{{1}}{{2}}}=\frac{1}{M+1}\big(l+\frac{1}{2}\big),\;l=0,\dots,M$. Compared to our previous work [38] that calculated spatial derivatives $\partial_{x}$ over a continuous domain, the spectral method results in a significantly more efficient expression for $\mathbf{f}(\mathbf{q})$ and allows to implement the nonlinear function (2) without approximation by Taylor series.
@@
$\mathbf{C}]_{ml}=\sqrt{\frac{2}{M+1}}\cos\bigg(\frac{\pi}{M+1}m\bigg(l+\frac{1}{2}\bigg)\bigg),\quad\begin{subarray}{c}m=1,\dots,M\\ l=0,\dots,M\end{subarray}.$
@@
$\mathbf{f}(\mathbf{q})=-\nabla_{\mathbf{q}}V(\mathbf{q}),\quad V(\mathbf{q})=\frac{1}{M+1}\sum\limits_{l=0}^{M}\mathcal{V}\big(\xi_{l+\nicefrac{{1}}{{2}}}\big).$ (8)
@@ 4.1  Quadratisation
where $\epsilon>0$ is a gauge constant. The auxiliary variable $\psi$ is well defined if the potential $V(\mathbf{q})$ (8) is non-negative which is the case for the chosen nonlinear function (2).
@@
$\begin{cases}\dot{\mathbf{q}}=\mathbf{p}\\ \dot{\mathbf{p}}=-2\bm{\Sigma}\mathbf{p}-\bm{\Omega}^{2}\mathbf{q}-\nu^{2}\mathbf{g}\psi+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}(t)\\ \dot{\psi}=\mathbf{g}^{\mathrm{T}}\mathbf{p}\end{cases}$ (9)
@@
Furthermore, we add a control term to the system (9) as described by Risse et al. [32] to reduce the drift between numerical values of $\psi$ and $\sqrt{2V(\mathbf{q})+\epsilon}$ in discrete time. This results in modification of the coupling term $\mathbf{g}$:
@@ 4.2  Time Discretisation
$\begin{cases}\delta_{t+}\mathbf{q}^{n-\nicefrac{{1}}{{2}}}=\mathbf{p}^{n}\\ \begin{split}\delta_{t+}\mathbf{p}^{n}=&-2\bm{\Sigma}\mu_{t+}\mathbf{p}^{n}-\bm{\Omega}^{2}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}-{}\\ &{}-\nu^{2}\mathbf{g}^{n}\mu_{t+}\psi^{n}+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}\end{split}\\ \delta_{t+}\psi^{n}=(\mathbf{g}^{n})^{\mathrm{T}}\mu_{t+}\mathbf{p}^{n}\end{cases}$ (10)
@@
Numerical energy for the scheme (10) is conserved and takes the form:
@@
Defining $\mathbf{q}^{n}\triangleq\mu_{t+}\mathbf{q}^{n-\nicefrac{{1}}{{2}}}$, the scheme (10) can be written in an update form $(\mathbf{q}^{n},\mathbf{p}^{n},\psi^{n})\rightarrow(\mathbf{q}^{n+1},\mathbf{p}^{n+1},\psi^{n+1})$:
$\begin{cases}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}=\mathbf{q}^{n}+\tfrac{k}{2}\mathbf{p}^{n}\\ \begin{split}\mathbf{p}^{n+1}&=\Big[\mathbf{I}+k\bm{\Sigma}+\tfrac{k^{2}\nu^{2}}{4}\mathbf{g}^{n}(\mathbf{g}^{n})^{\mathrm{T}}\Big]^{-1}\times{}\\ &{}\times\Big[\Big(\mathbf{I}-k\bm{\Sigma}-\tfrac{k^{2}\nu^{2}}{4}\mathbf{g}^{n}(\mathbf{g}^{n})^{\mathrm{T}}\Big)\mathbf{p}^{n}+{}\\ &{}+k\Big(-\bm{\Omega}^{2}\mathbf{q}^{n+\nicefrac{{1}}{{2}}}-\nu^{2}\mathbf{g}^{n}\psi^{n}+\bm{\Phi}(x_{\mathrm{e}})f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}\Big)\Big]\end{split}\\ \mathbf{q}^{n+1}=\mathbf{q}^{n+\nicefrac{{1}}{{2}}}+\tfrac{k}{2}\mathbf{p}^{n+1}\\ \psi^{n+1}=\psi^{n}+k(\mathbf{g}^{n})^{\mathrm{T}}\tfrac{\mathbf{p}^{n+1}+\mathbf{p}^{n}}{2}\end{cases}$ (11)
The inverse appearing in (11) can be easily computed with the Sherman-Morrison formula [35]. Using (5), we obtain an audio output as $w^{n}=\bm{\Phi}^{\mathrm{T}}(x_{\mathrm{o}})\mathbf{q}^{n}$.
@@ 5.1  Neural Ordinary Differential Equations
$\frac{d\mathbf{y}}{dt}=\mathbf{h}_{\theta}(\mathbf{y},t),\quad\mathbf{y}(0)=\mathbf{y}_{0}.$ (12)
Here $\mathbf{y}=\mathbf{y}(t)\!\!:\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is an unknown function of time $t$, $\mathbf{y}_{0}\in\mathbb{R}^{K}$ is an initial condition and $\mathbf{h}_{\theta}(\mathbf{y},t)\!\!:\mathbb{R}^{K}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{K}$ is a neural network where $\theta$ denotes the set of all learnable parameters and $K$ denotes the state dimension. Generally, a simple architecture such as an MLP is chosen for $\mathbf{h}_{\theta}(\mathbf{y},t)$. Chen et al. [6] have showed that the system (12) in combination with a numerical solver, labelled as ODENet, can be trained from the observed state data to reproduce the dynamics of a target system for which the theoretical model may be unknown.
Assume a target trajectory $\{\mathbf{y}_{0},\mathbf{y}^{1},\dots,\mathbf{y}^{N-1}\}$ sampled on the time grid $\{t^{n}\}_{n=0}^{N-1}$. Given a predicted trajectory $\{\mathbf{y}_{0},\tilde{\mathbf{y}}^{1},\dots,\tilde{\mathbf{y}}^{N-1}\}$ by a numerical solution to the initial value problem (12), i.e., a forward pass of the ODENet, we can construct an objective function $J(\theta)$ such as mean squared error (MSE):
$J(\theta)=\operatorname{MSE}(\tilde{\mathbf{y}}^{n},\mathbf{y}^{n})\triangleq\frac{1}{KN}\sum_{n=0}^{N-1}\left\lVert\tilde{\mathbf{y}}^{n}-\mathbf{y}^{n}\right\rVert_{2}^{2},$ (13)
@@ 5.2  Extension for Modal Synthesis
  • set $\bm{\Sigma},\bm{\Omega},\nu,\bm{\Phi}(x_{\mathrm{e}})$ in (14) using physical parameters of a target solution;
  • precompute $f_{\mathrm{e}}^{n+\nicefrac{{1}}{{2}}}=f_{\mathrm{e}}\big(t^{n+\nicefrac{{1}}{{2}}}\big),\;n=0,\dots,N-1$ using (3) and excitation parameters of a target solution;
  • use the numerical solver (11) as in the case of a regular system to produce a predicted trajectory.
The formulation (14) has strong implications. The exact expression for linear vibration exploits the periodic, harmonic and lossy nature of a musical system. Thus, we aid the optimisation process by constraining the space of possible solutions and improve interpretability of the model. Furthermore, the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ is memoryless and dimensionless, thus does not depend on physical parameters of a system and external excitation. Theoretically, these parameters can be changed after the training as long as range of displacements $\mathbf{q}$ stays the same as in a training dataset to simulate other configurations of a system. Some limitations exist since the neural network $\mathbf{f}_{\theta}(\mathbf{q})$ implicitly depends on boundary conditions and the number of modes.
@@ 5.3  Gradient Networks
$\mathbf{f}_{\theta}(\mathbf{q})=-\mathbf{W}^{\mathrm{T}}[\bm{\alpha}\odot\sigma(\mathbf{z})],\quad\mathbf{z}=\bm{\beta}\odot\mathbf{W}\mathbf{q}+\mathbf{b},$ (15)
@@
We can notice a strong resemblance between the GradNet (15) and the spectral method (Section 3.1). The weight matrix $\mathbf{W}$ acts as a transform to a hidden space where we apply the elementwise nonlinear function $\sigma(\mathbf{z})$ before performing an “inverse” transform with the transposed weight matrix $\mathbf{W}^{\mathrm{T}}$. Moreover, the potential (8) arising from the spectral method is expressed as a sum of convex ridge functions and the GradNet (15) with a monotonically-increasing activation function $\sigma$ and non-negative scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ can universally approximate gradients of such functions [5].
@@ 6  EVALUATION
For the parametrisation of $\mathbf{f}_{\theta}(\mathbf{q})$ we used the GradNet (15) with hidden dimension $H=1000$. This number was chosen based on available GPU memory since the validation loss was found to consistently decrease with increasing hidden dimension. However, smaller networks may also be sufficient from the perceptual point of view. A leaky rectified linear unit was used as an activation $\sigma$ which is a monotonically-increasing function with a non-negative antiderivative $\phi$. Kaiming initialisation [12] was used for the initial weights $\mathbf{W}$ and the biases $\mathbf{b}$ were initialised to zero. Logarithms of scaling vectors $\bm{\alpha}$ and $\bm{\beta}$ were used as learnable parameters to enforce non-negativity and were initialised around zero using a normal distribution.
@@
The training loss was the MSE (13) taken over the state vector $\mathbf{y}^{\mathrm{T}}=[\mathbf{q}^{\mathrm{T}},\mathbf{p}^{\mathrm{T}}]$, excluding the auxiliary variable $\psi$ arising from the SAV method. This implied that the potential $V_{\theta}(\mathbf{q})$ could be shifted by a constant relative to the target potential (8) since dynamics of the system and the training loss would not be affected. Backpropagation was performed using internal operations of the numerical solver (11), i.e., the “discretise-then-optimise” method, which is generally a preferred approach due to its gradient accuracy, speed and straightforward implementation [13]. Calculation of the control term $\mathbf{g}_{\mathrm{mod}}$ was excluded from the computational graph. In some cases, it was found to significantly compromise training by steering the potential $V_{\theta}(\mathbf{q})$ to incorrect values of the auxiliary variable $\psi$ and lead to an unexpected increase in the training loss. As only gradient calculation is affected, stability and drift regulation properties of the numerical solver (11) still hold during inference. The Adam optimiser [14] was used with default parameters. The training was performed for 2000 epochs using a cluster of four GPUs to parallelise batches of data. The resulting model was chosen based on the lowest validation loss obtained during optimisation.
@@
For training we used a variation of the teacher forcing technique [11] by splitting up a target trajectory into $\mathrm{1}$$\mathrm{ms}$ segments and providing true initial conditions for each segment to the ODENet. Since the numerical solver (11) is given as a one step update we had access to both displacement and velocity of the target numerical solution at each time step, and thus initial conditions for each segment. The auxiliary variable was initialised as $\psi_{0}=\sqrt{2V_{\theta}(\mathbf{q}_{0})+\epsilon}$ using initial displacements $\mathbf{q}_{0}$ to produce a consistent initial condition [32]. In addition, the excitation function (3) was shifted in time to reflect a new starting point for integration. The main reason for using this technique was to speed up training as the number of integration steps in the numerical solver is significantly reduced. These integration steps cannot be parallelised in time. Moreover, the likelihood of vanishing and exploding gradient problems during optimisation is also reduced by this technique as backpropagation on long time series is avoided [13].